OpenStudy (anonymous):

1-2cos^2x/1-2cosx sinx= sinx+cosx/sinx-cosx

OpenStudy (anonymous):

yes

hartnn (hartnn):

we have used \(\sin^2x+\cos^2x=1\) , so many times! but this is the type of problem when we use \(\Large 1 = \sin^2x=\cos^2x\)

hartnn (hartnn):

i mean \(\Large 1= \sin^2x+\cos^2x\)

hartnn (hartnn):

so, take left side and substitute this for 1 what u get ?

OpenStudy (anonymous):

sin^2x+cos^2x-2cos^2x/sin^2x+cos^2x-2cosxsinx

hartnn (hartnn):

simplify the numerator and notice that denominator is of the form a^2+b^2-2ab what does that equal ?

OpenStudy (anonymous):

I don't know, type the answer

hartnn (hartnn):

oh, really ? sorry, this is not an answering site! its a learning site :) and try the numerator part , its easy!

OpenStudy (anonymous):

ok sinx + cosx - 2cos^2x

OpenStudy (anonymous):

How do I simplify

hartnn (hartnn):

the num is actually sin^2x+cos^2x-2cos^2x right ? whats cos^2x-2cos^2x =.....? factor out cos^2 x

OpenStudy (anonymous):

Can you cross out the cos^2x and be left with sin^2x -2

hartnn (hartnn):

not actually, when we factor out, cos^2x-2cos^2x = cos^2x (1-2) = - cos^2 x got this ?

OpenStudy (anonymous):

ok is the numerator sin^2x + sin^2x. I factored out a cos^2x, then I had 1-cos^2x, which is sin^2x

hartnn (hartnn):

we were talking about num. only, which came out to be sin^2x -cos^2x isn't it ?

hartnn (hartnn):

sin^2x + cos^2x-2cos^2x =sin^2x + cos^2x (1-2) = sin^2 x- cos^2 x like this.

OpenStudy (anonymous):

I don't know im so lost right now

hartnn (hartnn):

look over the steps i gave you and ask in which step you have doubt

OpenStudy (anonymous):

ok where does the (1-2) come from? How did you get that

hartnn (hartnn):

you know factoring ? like ab+ac = a(b+c) <<<<<[factoring out a] so, cos^2x-2cos^2x = cos^2x (1-2) <<<<[factoring out cos^2x]

OpenStudy (anonymous):

Can it be written as 1-cos^2x though

hartnn (hartnn):

nopes, we need to factor OUT cos^2x from BOTh therms

OpenStudy (anonymous):

Oh ok now I see it

OpenStudy (anonymous):

I got that part

OpenStudy (anonymous):

then what?

hartnn (hartnn):

now a^2-b^2 = (a+b)(a-b) so, what about cos^2x-sin^2x = .... ?

OpenStudy (anonymous):

(cosx +sinx)(cosx-sinx)

OpenStudy (anonymous):

But its sin^2x + cos^2x, not -

hartnn (hartnn):

it is - sin^2x + cos^2x-2cos^2x =sin^2x + cos^2x (1-2) = sin^2 x- cos^2 x like this.

OpenStudy (anonymous):

So you multiplied (1-2) through that

OpenStudy (anonymous):

duh (1-2) is -1, then you times by a negative, which change the sign

hartnn (hartnn):

yes, thats correct. glad you understood by yourself.

OpenStudy (anonymous):

now the denominator

hartnn (hartnn):

the denom is of the form a^2+b^2-2ab which is = (a-b)^2

OpenStudy (anonymous):

which would be (sinx - cosx)^2

hartnn (hartnn):

absolutely correct! and what gets cancelled out from numerator and denominator ?

OpenStudy (anonymous):

The whole thing, top and bottom are both the same

hartnn (hartnn):

not actually

OpenStudy (anonymous):

No wait, (sinx - cos)^2 = (sinx-cos) (sinx-cosx), right?

hartnn (hartnn):

correct!

hartnn (hartnn):

what gets cancelled ? and what remains ?

OpenStudy (anonymous):

sinx + cosx over sinx - cosx, which equals the right side

OpenStudy (anonymous):

Omg your genius, thank you very much

hartnn (hartnn):

you are welcome ^_^